solve equation
x-3/5=x-3/x

Answer by kiru_khandelwal(79)

x - 3/5 = x - 3/x
=> (5x-3)/5 = (x^2 - 3)/x
=> (5x-3)*x = (x^2 - 3)*5
=> 5x^2 - 3x = 5x^2 -15
=> 5x^2 - 5x^2 -3x = -15
=> -3x = -15
=> 3x = 15
=> x = 15/3 = 5